If a nine or an eleven occurs on the first roll, the player wins. What is the probability of winning on the first roll?
Answer by jim_thompson5910(35256)

I'm assuming you're rolling two dice.

There are 36 possible outcomes (6*6 = 36).

There are 4 ways to roll a 9 (3+6=9,4+5=9,5+4=9,6+3=9)

There are 2 ways to roll a 11 (5+6=11,6+5=11)

There are 2+4 = 6 ways to roll either a 9 or 11.

The probability of winning on the first roll is 6/36 = 1/6